Q.2 The following equations describe an economy:

C= 10 + 0.5 Y (Consumption function)

I = 190-20i (Investment function)

Derive the equations for IS curve and represent it graphically for i=2 and i=5


1
Expert's answer
2021-05-27T11:03:06-0400

"Y=C+I"

"Y=10+0.5Y+190-20i"

"Y-0.5Y=200-20i"

"0.5Y=200-20i"

"Y= 400-40i"

The above is IS equation.

When "i=2," then output="400-40(2)"

"=400-80 = 320"

When "i=5," then output "=400-40(5)"

"=400-200 =200"

Its graphical representation is shown below-



The above figure shows there is an inverse relationship between ROI and output level. When i=2% then Y=320 and when ROI increases to 5% then output falls to 200.